# Adversarial Pruning: A Survey and Benchmark of Pruning Methods for Adversarial Robustness

### v1

@@ 6 Conclusion and Future Work
Adversarial Pruning (AP) methods represent an important set of techniques to reduce a model’s size while preserving robustness against adversarial attacks. In this regard, it is important, given their complexity and diversity, to have a systematic classification and analysis of the methods and their designs. In this work, we proposed a taxonomy for AP methods, with the goal of comparing the current and future AP methods’ designs. In addition, we reviewed existing AP’s adversarial evaluations and found multiple issues related to their reliability. Towards a fair, reliable comparison of AP methods, we thus proposed our benchmark, which allows AP methods to be evaluated in the same experimental and adversarial setup. Then, we used our benchmark to re-evaluate the available implementation of AP methods, which we studied alongside the taxonomy to associate the different designs with their effects. With our work, we aim to help the design of AP methods, which represent a highly powerful technique to guarantee both the efficiency and robustness of ML models. For future directions, we reckon that an ablation analysis might be ideal to study the effect of choosing different pipelines and specifics, as well as the extension to multiple models and datasets not limited to vision applications. In addition, we plan to extend our benchmark to multiple threat models and datasets.
@@ Acknowledgements
This work was carried out while G. Piras was enrolled in the Italian National Doctorate on AI run by the Sapienza University of Rome in collaboration with the University of Cagliari. This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU - NGEU; by the European Union’s Horizon Europe Research and Innovation Programme under the project Sec4AI4Sec, grant agreement No 101120393; and by Fondazione di Sardegna under the project “TrustML: Towards Machine Learning that Humans Can Trust”, CUP: F73C22001320007.

### v2

@@ 6 Conclusion and Future Work
In this work, we proposed a taxonomy of AP methods with the goal of comparing the current AP methods’ designs. In addition, we reviewed existing AP’s adversarial evaluations and found multiple issues related to their reliability. Towards a fair, reliable comparison of AP methods, we thus proposed our benchmark, which allows AP methods to be evaluated in the same experimental and adversarial setup. Hence, we used our benchmark to re-evaluate the available implementation of AP methods, which we studied alongside the taxonomy to associate the different designs with their effects. With our work, we provide a unique framework to “unwrap” and explain the design of these methods, while also allowing to test them uniformly and compare the results. The ultimate goal is thus to pave the way for future adversarial pruning research, which can now benefit from a reference point on which to compare and test their novel designs.
In conclusion, we acknowledge that our work is limited to neural networks, not expanding on more complex architectures such as Vision Transformers (ViT). However, unfortunately, work on ViT robustness and ViT pruning follow two different paths which have yet to meet. While existing work on pruning ViT [71] provide the foundation in fact, integrating these methods with adversarial robustness represents a non-trivial and yet unexplored challenge. In this regard, we believe that our taxonomy and benchmark can be leveraged by future work to investigate AP on transformers. Additionally, different applications of adversarial pruning, such as in natural language processing, remain likewise rather unexplored. We thus hope that this work inspires future research to tackle these challenges and advance adversarial pruning toward broader applicability and impact.
@@ Acknowledgements
This work was carried out while Giorgio Piras was enrolled in the Italian National Doctorate on AI run by the Sapienza University of Rome in collaboration with the University of Cagliari. This work was partially supported by the NRRP MUR program funded by the EU-NGEU under the projects SERICS (PE00000014) and FAIR (PE00000013); by the European Union’s Horizon Europe Research and Innovation Programme under the project Sec4AI4Sec (grant agreement no. 101120393) and ELSA (grant agreement no. 101070617); and by Fondazione di Sardegna under the project “TrustML: Towards Machine Learning that Humans Can Trust”, CUP: F73C22001320007.